OpenStudy (anonymous):

Trigonometry find positive angle less than 360 degrees or 2π that is coterminal with the given angle Problem: 23π/5

OpenStudy (anonymous):

Fist divided by \(2\pi\), what do you get?

OpenStudy (anonymous):

i did 23π-10π

OpenStudy (mertsj):

\[\frac{23\pi}{5}=4 \frac{3}{5}\pi-4\pi=\frac{3}{5}\]

OpenStudy (mertsj):

3/5 pi

OpenStudy (anonymous):

ok got the first part, how did you get 4π?

OpenStudy (mertsj):

Each revolution is 2pi so, to get rid of the maximum number of complete revolutions contained in 4 3/5 pi, I subtracted 4pi which is a multiple of 2 pi

OpenStudy (anonymous):

ok i understand that now. thank you

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Mertsj can you check your inbox